Circular No. 8772

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 M4 (SWAN) C. E. Woodward and M. S. Kelley, University of Minnesota; and

P. M. Hinz, M. A. Kenworthy, and W. F. Hoffman, Steward Observatory, University of Arizona, report the discovery, in 10.55-μm narrow-band (FWHM = 0.97 μm) images obtained with the 6.5-m MMT telescope (+ MIRAC4-BLINC) and adaptive optics (AO) secondary on Nov. 7.08 UT, of a detached-coma feature located 3′′.4 from the nuclear condensation of comet C/2006 M4 (corresponding to 2700 km at a geocentric distance of 1.10 AU) in p.a. ≈ 43^{o}. Aperture photometry, derived from images obtained through thin cirrus, with the AO system locked on the diffuse comet nucleus in a 3′′.1 circular beam, yields 10.55-μm fluxes of 9.5 ± 0.5 Jy for the nuclear condensation and 4.1 ± 0.4 Jy for the detached-coma feature. The detached-coma feature is elongated along the same position angle, with a spatial extent of ≈ 3′′ by 1′′.5. The nucleus was unresolved in the diffractionlimited images (FWHM = 0′′.4) and appears to be embedded in a diffuse coma (FWHM = 0′′.7). Further mid-infrared observations are encouraged.

COMETS C/2006 R3, C/2006 R4, C/2006 S7–S9 (SOHO) Several additional apparent comets have been found on SOHO website

images (cf. IAUC 8740), all being Kreutz sungrazers except for C/2006 R3 (Meyer group), which was very small, stellar in appearance, elongated, and of mag ∼ 7.5. K. Battams notes that C/2006 R4 was mostly stellar in appearance with a hint of tail, reaching mag ∼ 3.5. C/2006 S7 shows a hint of tail and was mostly stellar in appearance (mag ∼ 5) in C3 images, while in C2 images it appeared very diffuse and quite faint. C/2006 S8 was slightly elongated and reached mag ∼ 4.5 in C3 images; in C2 images, it was somewhat diffuse with a hint of tail, and it faded quite rapidly. C/2006 S9 appeared stellar with a hint of elongation (mag ∼ 6) in C3 images; in C2 images, it was diffuse with a hint of tail and quite faint.

Comet  2006 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2006 R3 Sept. 2.675 10^{h}42.^{m}6 + 9^{o}38′ C2  HS 2006-U63
C/2006 R4 11.746 10 53.9 + 3 18 C3 TH 2006-U63
C/2006 S7  18.138 11 20.8 + 0 52 C3/2 RM 2006-U63
C/2006 S8  18.488 11 22.4 − 0 03 C3/2 RM 2006-U63
C/2006 S9  18.488 11 27.3 + 0 54 C3/2 RM 2006-U63
2006 November 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green